# Revisiting electron-capture decay for Galactic cosmic-ray data

### v1

@@ 5 Conclusions
Last, we have compared the decayed fraction to both IS and TOA current data precision. While IS data are more favourable in terms of energy, only elemental fluxes – in which the isotope decay is diluted by the presence of the other stable isotopes – have been measured. Low-energy TOA data are more numerous and benefit from the latest generation of high-precision experiments (mainly ACE-CRIS for isotopes and fluxes, and AMS-02 for fluxes). All in all, only 49V, 51Cr and 57Co meet the EC-decay detectability condition – i.e., the predicted decayed fraction is larger than the current data precision –, in both isotopic TOA data and elemental IS or TOA fluxes. We stress that all our calculations were performed neglecting energy losses, so that the decayed fractions are probably slightly overestimated.
In summary, while the impact of EC decay remains difficult to confirm in current CR, the isotopes of V, Cr, and Co are the most promising candidates to do so. Several aspects of our calculation require refinement, though: first, the model must incorporate energy losses and solve the transport equation with the complete network of nuclear production cross-sections; second, uncertainties on nuclear cross-sections for $Z>30$ must be quantified, as they are known to bring large uncertainties in the modelling, especially for pure secondary species. In any case, the presence of many short-lived EC-unstable isotopes (with $Z>30$) that should be fully decayed, combined with the prospect of measurements in the near future (e.g., from ISS-TIGER), provides a strong motivation to investigate further this topic.
@@ Acknowledgments
This work is part of the project “Cosmic ray antideuterons as a probe for new physics” with project number OCENW.KLEIN.387 (Budget Number 11680) of the research programme Grant Open Competition Domain Science, which is financed by the Dutch Research Council (NWO).
@@ A.2 Stripping
where $C_{1}=0.285$ and $C_{2}=0.048$ are constants for capture from K-shell. The $Z_{\rm T}$ term arises from a sum over the contributions from all individual electrons in the target atom, while the $Z_{\rm T}^{2}$ term accounts for the ionisation by the nucleus. The final stripping cross-section formula is the weighted sum of the cross-sections for the different ISM target abundances.
@@ A.3 Extension for the multi-level model
All the formulas presented in the previous sections consider just two ionised states for the projectile, with effective charge $q_{\rm P}=Z_{\rm P}$ and $q_{\rm P}=Z_{\rm P}-1$. Indeed, they were specifically derived for attachment of an electron to a fully ionised nucleus or stripping from a hydrogenic atom. In our analysis, we generalise the transport equation to a multi-level model (see B.1), tracking the evolution of the CR ions up to any level of ionisation state $q_{\rm P}\in\{Z_{\rm P},(Z_{\rm P}-1),(Z_{\rm P}-2),\dots\}$. This requires the knowledge of the attachment and stripping cross-sections for partly ionised ions. The semi-empirical Schlachter formula proposes $\sigma^{\rm att}\propto q_{\rm P}^{3.9}$ (Schlachter et al., 1983), with an accuracy $\lesssim 2$ (Tolstikhina and Shevelko, 2023). As there are large uncertainties on this formula, we keep the description simple and assume the above formulae apply for all ionisation states. This assumption was actually made in the context of the more complex problem of the study of charge states of relativistic heavy ions in matter (Scheidenberger et al., 1998). In any case, we find that only attachment levels up to a few are relevant for most of the species in the CR context (see Sect. 3.4), so this approximation should not be too critical. Another assumption we make is to neglect multielectron capture or stripping – we assume electrons are attached or stripped one at a time –, whereas it can be significant at almost all energies (Tolstikhina and Shevelko, 2023). However, the dependence of the multielectron cross-sections on several parameters (atomic structure of colliding particles, velocity of the incident ion, and other atomic parameters) is still being investigated experimentally and theoretically (e.g., Wu et al., 2025). In the end, for the same reasons as above – simplicity and the fact that mostly low-level attachments and stripping seem to be relevant for CR –, this assumption should be good enough in the CR context.
@@ Appendix B Solutions of the transport equation
and
The (semi-) analytical solutions of the transport equation, for stable and $\beta$-radioactive GCR ions, (with) without energy redistribution, are gathered in Maurin (2020), for the LBM and 1D-DM. In this appendix, we extend these solutions to the EC case, without energy redistribution and without galactic wind (for the 1D-DM).
Compared to the standard solutions, we now have to follow the various ionisation states of the GCR species. Indeed, whereas CRs are expected to be fully ionised at high-enough energy (see Fig. 1 in Sect. 2.3), several ionisation states may co-exist at lower energy (Letaw et al., 1984b, 1985). As a result, instead of solving the transport equation for the differential density of the CR ion $N$, we have to solve it for a tower of ionised states, denoted below $N_{0}$ (fully ionised), $N_{1}$ (one $e^{-}$ attached), $N_{2}$ (two $e^{-}$ attached), etc. Each level is connected to the one below and above via the electron attachment and stripping rates, $\Gamma^{\rm att}$and $\Gamma^{\rm str}$, the decay rate $\Gamma^{\rm EC}=1/(\gamma\tau_{EC})$ being enabled for all $N_{j>0}$ (i.e., at least one electron attached). The total measured flux of this CR species is then given by
$N_{(n+1)\penalty 10000\ {\rm lev}}=N_{0}+N_{1}+N_{2}+\dots+N_{n}\,$, (19)
@@ B.1 Multi-level solution for EC decay
We consider below a single EC-unstable CR species, and we assume a generic source term $\cal S$} (of primary or secondary origin), providing the ionised state $N_{0}$ only. In this first step, we neglect energy redistribution terms, and for simplicity, we also discard the convective term in the 1D-DM. The multi-level (up to level $n$) transport equation for the coupled ionised states, for any of the above models, can be written generically (with $j=1,\dots,n-1$)
@@
We stress that the dimension of the latter quantity is a power-law of time T, $\theta_{j}]=$ Tj-1-n, so that we recover $N_{j}]=[Q]$ T-1.

### v2

@@ 5 Conclusions
Last, we have compared the decayed fraction to both IS and TOA current data precision. While IS data are more favourable in terms of energy, only elemental fluxes – in which the isotope decay is diluted by the presence of the other stable isotopes – have been measured. Low-energy TOA data are more numerous and benefit from the latest generation of high-precision experiments (mainly ACE-CRIS for isotopes and fluxes, and AMS-02 for fluxes). All in all, only 49V, 51Cr and 57Co meet the EC-decay detectability condition – i.e., the predicted decayed fraction is larger than the current data precision –, in both isotopic TOA data and elemental IS or TOA fluxes. Actually, the daughters of 49V, 51Cr and 55Fe, fed by their parent decays, are even more significantly impacted (than their parents): parent-to-daughter ratios (of these species) are the best hope for a detection at this stage. We stress that all our calculations were performed neglecting energy losses, so that the decayed fractions are probably slightly overestimated. This means that the results obtained are conservative with respect to the non-detectability cases, but detectable ones need to be confirmed by a more rigorous calculation.
In summary, while the impact of EC decay remains difficult to confirm in current CR data, the isotopes of V, Cr, and Co are the most promising candidates to do so. Several aspects of our calculation require refinement, though: first, the model must incorporate energy losses and solve the transport equation with the complete network of nuclear production cross-sections; second, uncertainties on nuclear cross-sections for $Z>30$ must be quantified, as they are known to bring large uncertainties in the modelling, especially for pure secondary species. In any case, the presence of many short-lived EC-unstable isotopes (with $Z>30$) that should be fully decayed, combined with the prospect of measurements in the near future (e.g., from ISS-TIGER), provides a strong motivation to investigate further this topic.
@@ Acknowledgments
We thank the referee for their insightful comments. This work is part of the project “Cosmic ray antideuterons as a probe for new physics” with project number OCENW.KLEIN.387 (Budget Number 11680) of the research programme Grant Open Competition Domain Science, which is financed by the Dutch Research Council (NWO). The work of D. Maurin was supported by the INTERCOS project funded by IN2P3.
@@ A.2 Stripping
where $C_{1}=0.285$ and $C_{2}=0.048$ are constants for capture from K-shell. The $Z_{\rm T}$ term arises from a sum over the contributions from all individual electrons in the target atom, while the $Z_{\rm T}^{2}$ term accounts for the ionisation by the nucleus. The final stripping cross-section formula is the weighted sum of the cross-sections for the different ISM target abundances. We consider here the contributions of H and He to the ISM target material, with with their number fractions being 0.9 and 0.1, respectively.
@@ A.3 Extension for the multi-level model
All the formulas presented in the previous sections consider just two ionised states for the projectile, with effective charge $q_{\rm P}=Z_{\rm P}$ and $q_{\rm P}=Z_{\rm P}-1$. Indeed, they were specifically derived for attachment of an electron to a fully ionised nucleus or stripping from a hydrogenic atom. In our analysis, we generalise the transport equation to a multi-level model (see B.1), tracking the evolution of the CR ions up to any level of ionisation state $q_{\rm P}\in\{Z_{\rm P},(Z_{\rm P}-1),(Z_{\rm P}-2),\dots\}$. This requires the knowledge of the attachment and stripping cross-sections for partly ionised ions. The semi-empirical Schlachter formula proposes $\sigma^{\rm att}\propto q_{\rm P}^{3.9}$ (Schlachter et al., 1983), with an accuracy $\lesssim 2$ (Tolstikhina and Shevelko, 2023). As there are large uncertainties on this formula, we keep the description simple and assume the above formulae apply for all ionisation states. This assumption was actually made in the context of the more complex problem of the study of charge states of relativistic heavy ions in matter (Scheidenberger et al., 1998). In any case, we find that only attachment levels up to a few are relevant for most of the species in the CR context (see Sect. 3.4), so this approximation should not be too critical. Another assumption we make is to neglect multielectron capture or stripping – we assume electrons are attached or stripped one at a time –, whereas mutielectron processes can be significant at almost all energies (Tolstikhina and Shevelko, 2023). However, the dependence of the multielectron cross-sections on several parameters (atomic structure of colliding particles, velocity of the incident ion, and other atomic parameters) is still being investigated experimentally and theoretically (e.g., Wu et al., 2025). In the end, for the same reasons as above – simplicity and the fact that mostly low-level attachments and stripping seem to be relevant for CR –, this assumption should be good enough in the CR context.
@@ B.1 Multi-level solution for EC decay
We consider below a single EC-unstable CR species, and we assume a generic source term $\cal S$} (of primary or secondary origin), providing the ionised state $N_{0}$ only. In this study, we neglect energy redistribution terms, and for simplicity, we also discard the convective term in the 1D-DM. The multi-level (up to level $n$) transport equation for the coupled ionised states, for any of the above models, can be written generically (with $j=1,\dots,n-1$)
@@
We stress that the dimension of $\theta_{j}$ is a power-law of time T, $\theta_{j}]=$ Tj-1-n, so that we recover $N_{j}]=[Q]$ T-1.